# Car-Following Models: A Multidisciplinary Review

### v4

@@ IV-A Applications
Analyzing car-following behavior and developing intelligent control algorithms have significant consequences for road safety and energy efficiency. The profound impacts of the different driving algorithms (e.g., strategic models, model-predictive control, and RL controller) and various connected environments have set the stage for a new era of traffic flow dynamics. Many applications based on car-following models have been undertaken, summarized with the the following topics.
@@
  • Topic 1: Traffic Flow Dynamics. Simulation analysis through car following models offer insights into emerging traffic flow characteristics under conventional or connected environment. CF models serve as the basis for exploring how connected and automated mobility (CAM) could enhance traffic dynamics and address both the opportunities and challenges presented by mixed traffic scenarios [181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192].
@@
  • Topic 5: Cyber-Physical Human Systems. The seamless integration of human elements, computer-based algorithms, and physical components is essential to transform transportation systems into Cyber-Physical Human Systems (CPHS). By leveraging car following models, researchers are able to develop analytical frameworks that not only enhance the performance of these vehicles but also improve the overall security of the transportation system. This integration and testing pave the way for the eventual realization of fully autonomous driving capabilities [217, 218, 219, 220, 221, 222, 223, 224].
@@
  • Topic 7: Trajectory Reconstruction. Sensor-based trajectory acquisition often encounters issues with noise and missing information. This can be due to various factors such as sensor limitations (fixed location or penetration rates), environmental conditions, and the inherent complexity of traffic environments.This is where car following models become crucial to estimate the missing information in vehicle trajectories with the partial data collected from these sensors. The complete and continuous vehicle trajectory is vital for various applications. [232, 233, 234, 235, 236, 237, 238, 239, 240, 241, 242, 243, 244].
IV-B Dataset
Detailed high-quality vehicle trajectory data are of great significance for the development of driving behavior models and vehicle control algorithms. With realistic microscopic longitudinal and lateral maneuvers, the valuable vehicle trajectory data lay the foundation for many scientific discoveries. The collection of realistic vehicle trajectory data is identified as important as modeling and understanding the interactions of travelers. There are two types of vehicle driving datasets that are often used as benchmarks. 1. Lagrangian-type data collected from mobile devices on floating vehicles. 2. Eulerian-type traffic-flow data that monitors the entire traffic flow from stationary devices. An integrated Benchmark [245] for Car-Following Behavior Modeling is extracted from five public driving datasets under the same format and criteria, which consists of more than 80K car following events. A summary table of datasets suitable for car following model research is presented in Table VI. It is important to note that while this table showcases some of the most pivotal datasets, it is by no means exhaustive. There are undoubtedly other datasets available that could provide valuable contributions to the field. The datasets highlighted in our summary are primarily focused on open street environments capturing captures the complexity and unpredictability of naturalistic driving scenarios as opposed to closed testing environments.
TABLE VI: Summary of Trajectory Dataset Collected from both Mobile and Fixed Sensors
@@ IV-B Dataset
The NGSIM datasets provide an unprecedented depth of trajectory data, which were captured at four locations: The US-101 and I-80 sites offer insights into freeway conditions, while Peachtree Street and Lankershim Boulevard provide data on urban street environments.
Each dataset is segmented into three 15 minute periods, covering approximately 500 meters (1,600 feet) in length, with over 11.8 million rows of trajectory data.
Zen Traffic Data [247]
2018
@@
Zen Dataset is characterized by its high traffic volume, two lanes in each direction, and a complex road layout featuring S-shaped curves and elevation changes.
3600 trajectories on 2km highway segment for 1 hour.
HighD [248]
2018
@@
The HighD dataset uses drone on German highways over various sections of highways.
110,500 vehicles, 28,000 vehicle miles traveled, and 147 hours
ACFR five roundabouts dataset [249]
2019
@@
This dataset is aimed at studying naturalistic driving behaviors at unsignalized intersections. The dataset offers insights into a wide range of driving behaviors, such as negligence, speeding, and assertiveness.
A dataset containing over 23,000 vehicle recordings at five different roundabouts, captured by a LiDAR-based detection and tracking system.
INTERA- -CTION [250]
2019
@@
The INTERACTION dataset stands out for complex and critical driving situations, often within unstructured settings lacking clear right-of-way rules, where driver behavior is significantly influenced by the actions of others.
Roundabout Scenarios (10,479 vehicle trajectories). Unsignalized intersection scenarios (14,867 trajectories). The merging and lane change scenarios (10,933 trajectories). The signalized intersection (3,775 trajectories).
inD [251]
2020
The inD (intersection Drone) dataset is a collection of naturalistic road user trajectories at urban intersections.
@@
More than 11,500 road users, including vehicles, bicyclists and pedestrians at intersections with 10 hours of data from four intersections
rounD [252]
2020
The rounD dataset is a collection of road user trajectories at roundabouts in Germany, captured using a drone camera.
@@
A total of 6 hours of recordings with more than 13,746 road users including cars, vans, trucks, buses, pedestrians, bicycles, and motorcycles.
I-24 Motion [253]
2020
@@ I-24 Motion [253]
I-24 Motion stands out for its large-scale data acquisition, capturing detailed vehicle trajectories across multiple lanes and over an extended period.
With 276 cameras that cover 4.2 miles of I-24, the data collected is intended to support a wide range of applications.
pNEUMA [254]
2020
Coordinated Drone video-based vehicle trajectory collected in downtown area in Athens, Greece during morning peak hour.
@@
The pNEUMA dataset encompasses close to half a million vehicle trajectories over a congested 1.3 km² area, featuring more than 100 km-lanes or road network.
Ubiquitous Traffic Eyes [255]
2021
The database encompasses a wide array of road conditions and traffic conditions, with manual corrections ensured 100% vehicle detection and tracking accuracy.
@@
The data segments vary in duration, ranging from 4 to 20 minutes. This dataset is characterized by its high precision, with time accuracy up to 0.1 seconds and position accuracy up to 0.01 meters.
HIGH-SIM [256]
2021
High-Granularity Highway Simulation (HIGH-SIM) consists of 2-hour 30fps vehicle trajectory dataset captured by three 8K cameras on the helicopter.
@@
Aerial videos at multiple successive locations of Interstate-75 over 2,438m highway segment, containing data of three lanes and one off-ramp from 4:15 p.m. to 6:15 p.m.
AUTOMATUM DATA [257]
2021
AUTOMATUM Drone-based dataset includes a validation mechanism where trajectories of objects are compared against test vehicles.
@@
Generated from 30 hours video with 18,724 vehicle miles traveled, where the relative speed error is less than 0.2 percent
MAGIC [258]
2022
Coordinated Drone video-based vehicle trajectory collected in Shanghai, China during morning peak hour
@@
3 hours of aerial video at six points on 4,000 m of urban highway (2,000 m in each direction) with four main lanes, three off-ramps, and three on-ramps.
CitySim [259]
2022
@@
CitySim is Drone video-based vehicle trajectory data for safety-oriented research, providing seven bounding box points for fine-grained analysis, with the capability to facilitate digital-twin-related research.
The CitySim dataset provides more than 1,200 minutes of drone video data from multiple countries in diverse weather conditions and locations.
Safety Pilot Model Deployment (SPMD) [260]
2012
The SPMD program contains Basic safety messages (BSM), vehicle trajectory, driver-vehicle interaction, and weather data are collected to evaluate V2V communication.
@@
The SPMD data sets contain sanitized mobility data elements collected from about 3,000 vehicles, equipped with connected vehicle technologies, traversing Ann Arbor, Michigan.
Naturalistic Driving Study (NDS) [261]
2013
@@
This dataset was collected through an extensive observational study with advanced data acquisition systems, including CAN-Bus, GPS, outward and inward facing cameras, and forward radar to capture driver behavior, vehicle movement, and environmental conditions.
More than 5 million trip files from nearly 3500 drivers located in six geographic regions across the United States.
nuScenes [262]
2020
The nuScenes dataset incorporates 6 cameras, 5 radars, and 1 LiDAR sensor. The nuScenes dataset is notable for its holistic view of detailed maps, weather conditions, and day/night cycles.
@@
With almost 1.4 million 3D bounding boxes across 23 object classes and over 8 million annotated instances, the nuScenes dataset spans over 100 hours of driving data, capturing various urban driving conditions.
Waymo Open Dataset [263]
2020
@@
The Waymo Open Dataset features millions of 3D bounding boxes for vehicles, pedestrians, cyclists, and other objects encountered in urban and suburban environments.
2,030 scenes for sensor data, and 103,354 scenes for motion data
Lyft Level 5 Dataset [264]
2021
@@
This dataset captured with a fleet of autonomous vehicles navigating through urban and suburban areas over a four-month period, enhanced with high-definition maps and traffic light statuses.
This dataset, consisting of over 170,000 scenes each lasting about 25 seconds, offering detailed perceptions and motion data.
V CONCLUSION AND OUTLOOK
@@
This review aims to enhance understanding of the underlying design principles in different categories of car-following (CF) models, enabling researchers to formulate problems with a suitable framework. As the path to achieve level-5 autonomy continues, incorporating general traffic flow knowledge presents both challenges and opportunities. The future direction of driver behavior models, particularly in the context of integrating Artificial General Intelligent (AGI) capabilities and learning efficiencies into car following models, presents a fascinating and complex challenge. Bridging the gap between conventional fixed driver behavior models and rapidly advancing AI models requires a multifaceted approach.
Hybrid-learning Systems: The future of car-following models likely lies in a synergistic combination of theory-based and data-driven approaches. One of hybrid-learning approach is Physics-informed machine learning (PIML), which offers a promising approach to enhance car following models. PIML allows for the integration of these established kinematic models with data-driven insights derived from real-world driving data. Another promising avenue for developing hybrid learning systems is Neural symbolic AI, which merges the pattern recognition capabilities of deep learning with the structured, rule-based logic of symbolic reasoning. This approach enables the creation of models that can not only learn from data, but also incorporate and respect the underlying physical laws that govern real-world phenomena. This hybrid approach can potentially create systems with built-in physical constraints, enhancing their robustness, interpretability and controllability.
World Model Predictive Control: Driving agent is expected to adapt to new environments, learn new tasks as fast as humans, or have a near-human level of intelligence. Therefore, model predictive control for next-generation car following should be based on common sense after harvesting enormous amounts of background knowledge, for instance, the ChatGPT-like world model. These Large Language Models (LLM) can make significant advancements in how these systems understand and react to complex driving scenarios. In contrast to the stimulus-response or feedback/feedforward control, intelligent driving agents should have internalized the world model utilizing the knowledge in the past to react to situations in the present, or before they arise in the future.
Hierarchical Planning and Control: One of the main features of human intelligence is that people can learn abstract concepts. Human perception systems process information about the surrounding environment in a hierarchical manner. Hierarchical Learning and planning capabilities enable the intelligent driving agent to capture dependencies of inputs based on abstract concepts with larger receptive fields. Many other factors, including vehicle, road, environment, traffic situation, etc. influence driver behavior characteristics. Hierarchical learning can encode driving characteristics and traffic environments into high-level representations to make complex action plans.
Multimodal Signals: Current CF models are fed with curated trajectory data after post-processing, aggregation, and sensor fusion. Future opportunities for building the CF model include incorporating the vision pipeline in training connected and automated vehicles. As for humans, much knowledge about the law of motion is learnable through pure observation. The human decision-making process is based on various senses. Therefore, the intelligent driving agent is expected to interpret information from multiple sources (Camera, LiDAR, Radar, Can-bus, Gyroscope, Psychology Signal, and Audio). The most common approaches are multimodal fusion, cross-modality, and shared representation learning.
Self-Supervised Learning: The Majority of existing CF models have specific objectives and rely on predefined parameters and labeled data. Self-Supervised Learning employs an energy-based model to estimate the energy of sequences of actions defined by compatibility. The vehicle trajectory plan is considered a temporal prediction scenario. Self-supervised learning captures the dependencies between past and future outputs through contrastive or regularized learning.
Exploring these directions requires interdisciplinary collaboration, combining insights from artificial intelligence, cognitive science, psychology, computer science, traffic engineering and automotive engineering.
VI acknowledgement

### v5

@@ V-B Applications
Analyzing car-following behavior and developing intelligent control algorithms have significant consequences for road safety and energy efficiency. The profound impacts of the different driving algorithms (e.g., strategic models, model-predictive control, RL controller, and LLM-agent) and various connected environments have set the stage for a new era of traffic flow dynamics. Many applications based on car-following models have been undertaken, summarized with the following topics.
@@
  • Topic 1: Traffic Flow Dynamics. Simulation analysis through car following models offer insights into emerging traffic flow characteristics under conventional or connected environment. CF models serve as the basis for exploring how connected and automated mobility (CAM) could enhance traffic dynamics and address both the opportunities and challenges presented by mixed traffic scenarios [211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223].
@@
  • Topic 5: Cyber-Physical Human Systems. The seamless integration of human elements, computer-based algorithms, and physical components is essential to transform transportation systems into Cyber-Physical Human Systems (CPHS). By leveraging car following models, researchers are able to develop analytical frameworks that not only enhance the performance of these vehicles but also improve the overall security of the transportation system. This integration and testing pave the way for the eventual realization of fully autonomous driving capabilities [248, 249, 250, 251, 252, 253, 254, 255, 256].
@@
  • Topic 7: Trajectory Reconstruction. Sensor-based trajectory acquisition often encounters issues with noise and missing information. This can be due to various factors such as sensor limitations (fixed location or penetration rates), environmental conditions, and the inherent complexity of traffic environments. This is where car following models become crucial to estimate the missing information in vehicle trajectories with the partial data collected from these sensors. The complete and continuous vehicle trajectory is vital for various applications. [264, 265, 266, 267, 268, 269, 270, 271, 272, 273, 274, 275].
V-C Dataset
Detailed high-quality vehicle trajectory data are of great significance for the development of driving behavior models and vehicle control algorithms. With realistic microscopic longitudinal and lateral maneuvers, the valuable vehicle trajectory data lay the foundation for many scientific discoveries. The collection of realistic vehicle trajectory data is identified as important as modeling and understanding the interactions of travelers. There are two types of vehicle driving datasets that are often used as benchmarks. 1. Lagrangian-type data collected from mobile devices on floating vehicles. 2. Eulerian-type traffic-flow data that monitors the entire traffic flow from stationary devices. An integrated Benchmark [276] for Car-Following Behavior Modeling is extracted from five public driving datasets under the same format and criteria, which consists of more than 80K car following events. A summary table of datasets suitable for car following model research is presented in Table VII. It is important to note that while this table showcases some of the most pivotal datasets, it is by no means exhaustive. There are undoubtedly other datasets available that could provide valuable contributions to the field. The datasets highlighted in our summary are primarily focused on open street environments capturing captures the complexity and unpredictability of naturalistic driving scenarios as opposed to closed testing environments.
TABLE VII: Summary of Trajectory Dataset Collected from both Mobile and Fixed Sensors
@@ V-C Dataset
Next Generation Simulation (NGSIM) [277] 2006 The NGSIM datasets provide an unprecedented depth of trajectory data, which were captured at four locations: The US-101 and I-80 sites offer insights into freeway conditions, while Peachtree Street and Lankershim Boulevard provide data on urban street environments. Each dataset is segmented into three 15 minute periods, covering approximately 500 meters (1,600 feet) in length, with over 11.8 million rows of trajectory data.
@@
Zen Traffic Data [278] 2018 Zen Dataset is characterized by its high traffic volume, two lanes in each direction, and a complex road layout featuring S-shaped curves and elevation changes. 3600 trajectories on 2km highway segment for 1 hour.
@@
HighD [279] 2018 The HighD dataset uses drone on German highways over various sections of highways. 110,500 vehicles, 28,000 vehicle miles traveled, and 147 hours
@@
ACFR five roundabouts dataset [280] 2019 This dataset is aimed at studying naturalistic driving behaviors at unsignalized intersections. The dataset offers insights into a wide range of driving behaviors, such as negligence, speeding, and assertiveness. A dataset containing over 23,000 vehicle recordings at five different roundabouts, captured by a LiDAR-based detection and tracking system.
@@
INTERA- -CTION [281] 2019 The INTERACTION dataset stands out for complex and critical driving situations, often within unstructured settings lacking clear right-of-way rules, where driver behavior is significantly influenced by the actions of others. Roundabout Scenarios (10,479 vehicle trajectories). Unsignalized intersection scenarios (14,867 trajectories). The merging and lane change scenarios (10,933 trajectories). The signalized intersection (3,775 trajectories).
@@
inD [282] 2020 The inD (intersection Drone) dataset is a collection of naturalistic road user trajectories at urban intersections. More than 11,500 road users, including vehicles, bicyclists and pedestrians at intersections with 10 hours of data from four intersections
@@
rounD [283] 2020 The rounD dataset is a collection of road user trajectories at roundabouts in Germany, captured using a drone camera. A total of 6 hours of recordings with more than 13,746 road users including cars, vans, trucks, buses, pedestrians, bicycles, and motorcycles.
@@
I-24 Motion [284] 2020 I-24 Motion stands out for its large-scale data acquisition, capturing detailed vehicle trajectories across multiple lanes and over an extended period. With 276 cameras that cover 4.2 miles of I-24, the data collected is intended to support a wide range of applications.
@@
pNEUMA [285] 2020 Coordinated Drone video-based vehicle trajectory collected in downtown area in Athens, Greece during morning peak hour. The pNEUMA dataset encompasses close to half a million vehicle trajectories over a congested 1.3 km² area, featuring more than 100 km-lanes or road network.
@@
Ubiquitous Traffic Eyes [286] 2021 The database encompasses a wide array of road conditions and traffic conditions, with manual corrections ensured 100% vehicle detection and tracking accuracy. The data segments vary in duration, ranging from 4 to 20 minutes. This dataset is characterized by its high precision, with time accuracy up to 0.1 seconds and position accuracy up to 0.01 meters.
@@
HIGH-SIM [287] 2021 High-Granularity Highway Simulation (HIGH-SIM) consists of 2-hour 30fps vehicle trajectory dataset captured by three 8K cameras on the helicopter. Aerial videos at multiple successive locations of Interstate-75 over 2,438m highway segment, containing data of three lanes and one off-ramp from 4:15 p.m. to 6:15 p.m.
@@
AUTOMATUM DATA [288] 2021 AUTOMATUM Drone-based dataset includes a validation mechanism where trajectories of objects are compared against test vehicles. Generated from 30 hours video with 18,724 vehicle miles traveled, where the relative speed error is less than 0.2 percent
@@
MAGIC [289] 2022 Coordinated Drone video-based vehicle trajectory collected in Shanghai, China during morning peak hour 3 hours of aerial video at six points on 4,000 m of urban highway (2,000 m in each direction) with four main lanes, three off-ramps, and three on-ramps.
@@
CitySim [290] 2022 CitySim is Drone video-based vehicle trajectory data for safety-oriented research, providing seven bounding box points for fine-grained analysis, with the capability to facilitate digital-twin-related research. The CitySim dataset provides more than 1,200 minutes of drone video data from multiple countries in diverse weather conditions and locations.
@@
Safety Pilot Model Deployment (SPMD) [291] 2012 The SPMD program contains Basic safety messages (BSM), vehicle trajectory, driver-vehicle interaction, and weather data are collected to evaluate V2V communication. The SPMD data sets contain sanitized mobility data elements collected from about 3,000 vehicles, equipped with connected vehicle technologies, traversing Ann Arbor, Michigan.
@@
Naturalistic Driving Study (NDS) [292] 2013 This dataset was collected through an extensive observational study with advanced data acquisition systems, including CAN-Bus, GPS, outward and inward facing cameras, and forward radar to capture driver behavior, vehicle movement, and environmental conditions. More than 5 million trip files from nearly 3500 drivers located in six geographic regions across the United States.
@@
nuScenes [293] 2020 The nuScenes dataset incorporates 6 cameras, 5 radars, and 1 LiDAR sensor. The nuScenes dataset is notable for its holistic view of detailed maps, weather conditions, and day/night cycles. With almost 1.4 million 3D bounding boxes across 23 object classes and over 8 million annotated instances, the nuScenes dataset spans over 100 hours of driving data, capturing various urban driving conditions.
@@
Waymo Open Dataset [294] 2020 The Waymo Open Dataset features millions of 3D bounding boxes for vehicles, pedestrians, cyclists, and other objects encountered in urban and suburban environments. 2,030 scenes for sensor data, and 103,354 scenes for motion data
@@
Lyft Level 5 Dataset [295] 2021 This dataset captured with a fleet of autonomous vehicles navigating through urban and suburban areas over a four-month period, enhanced with high-definition maps and traffic light statuses. This dataset, consisting of over 170,000 scenes each lasting about 25 seconds, offering detailed perceptions and motion data.
OpenACC [296] 2021 OpenACC dataset is collected for understanding of commercial ACC control logic and is designed for extensive research into vehicle performance and safety. The dataset is a collection from 27 vehicle models across 17 brands, featuring a mix of engine types and captured using advanced GPS systems under various driving conditions.
VI CONCLUSION AND OUTLOOK
@@
This review aims to enhance understanding of the underlying design principles in different categories of CF Models, enabling researchers to formulate problems with a suitable framework. As the path to achieve level-5 autonomy continues, incorporating general traffic flow knowledge presents both challenges and opportunities. The future direction of driver behavior models, particularly in the context of integrating human-like reasoning capabilities and learning efficiencies into car following models, presents a fascinating and complex challenge. Bridging the gap between conventional fixed and myopic driver behavior models and rapidly advancing AI models requires a multifaceted approach.
Hybrid-learning Systems: The potential of hybrid-learning systems lies in their ability to produce models that are not only highly predictive but also deeply understandable and controllable. Among the leading methodologies in this frontier is Physics-informed Machine Learning (PIML), which stands out for its ability to harmonize traditional kinematic models with the implicit understanding gleaned from vast datasets of real-world driving behaviors. PIML represents a significant step forward by embedding physical principles directly into machine learning frameworks, thereby ensuring that predictions not only fit observed data but also adhere to fundamental laws of motion. This alignment guarantees that the models remain valid under a wide range of conditions.
Another development in hybrid learning systems is Neural Symbolic AI, which seeks to bridge the gap between the intuitive pattern recognition capabilities of deep learning and the logical rigor of symbolic reasoning. By combining these two approaches, Neural Symbolic AI facilitates the construction of models that are both data-efficient and interpretable. These models excel in learning from available data while simultaneously encoding the structured knowledge that characterizes human reasoning processes. Such a dual approach enables the models to make sense of complex scenarios, predict outcomes, and make decisions that are both logically sound and grounded in real-world data. The inherent flexibility of Neural Symbolic AI allows it to adapt to new information and evolving scenarios, making it an ideal candidate for tackling the dynamic challenges of car-following behavior. As the field of autonomous driving continues to evolve, hybrid-learning systems stand poised to play a crucial role in developing models that seamlessly blend empirical data with theoretical insights, paving the way for the next generation of intelligent transportation systems that are both robust and adaptable.
World Model Predictive Control: The aspiration for driving agents to adapt swiftly to new environments and assimilate new tasks with near-human intelligence sets a high bar for the evolution of autonomous driving technologies. At the core of this evolution lies the concept of World Model Predictive Control, a paradigm shift that integrates common sense and a vast repository of background knowledge into the decision-making processes of next-generation intelligent driver models. This approach, inspired by advancements in Large Language Models (LLMs) like ChatGPT, represents a leap forward in enabling driving agents to comprehend and navigate complex driving scenarios with unprecedented depth and nuance.
The principle behind World Model Predictive Control is to endow driving agents with an internalized representation of the world. This internal model, enriched by extensive learning from diverse data sources, allows the agent to predict and plan several steps ahead, moving beyond reactive stimulus-response mechanisms to proactive situational management. Unlike traditional models that rely on immediate feedback or predefined control laws, this approach leverages predictive analytics to foresee and navigate future states, drawing on a comprehensive understanding of the driving environment’s intricacies. The integration of such a world model into predictive control systems harness the full power of artificial intelligence to ensure safety and efficiency on the roads.
Hierarchical Planning and Control: The ability to understand and apply abstract concepts is a defining feature of human intelligence. These frameworks structure decision-making across multiple layers, ranging from immediate sensory responses at the lowest level to strategic planning at the highest. Incorporating hierarchical learning into intelligent driving systems facilitates the encoding of diverse factors—such as vehicle characteristics, road conditions, and traffic dynamics—into high-level representations. This allows for the formulation of complex action plans that dynamically adapt to changing conditions, thereby enhancing operational safety and predictive accuracy. The intermediate levels of this hierarchy are particularly crucial for understanding and predicting the behaviors of nearby vehicles, improving the system’s ability to anticipate and react to potential hazards before they arise.
The integration of hierarchical planning and control into car-following models marks a significant step towards achieving true autonomy in driving technologies. By enabling vehicles to not only react to their immediate environment but also engage in anticipatory actions based on scenario analysis and predictive modeling, these systems enable agents to make decisions based on a blend of immediate data and abstract concepts. The development of such advanced models paves the way for a future where autonomous vehicles operate with a level of intuition and foresight akin to human drivers, contributing to safer, more efficient, and harmonious traffic ecosystems.
Multimodal Signals: The evolution of car-following models is increasingly pivoting towards the integration of multimodal signals to closely mimic the human driving experience, which inherently relies on a symphony of sensory inputs for decision-making. Current car-following models primarily utilize curated trajectory data, derived from post-processing, aggregation, and sensor fusion techniques. Human drivers base their decisions on a complex blend of visual cues, auditory signals, and even haptic feedback, something that the next generation of intelligent driving agents must emulate. By integrating information from a diverse array of sources—such as Cameras, LiDAR, Radar, Can-bus systems, Gyroscopes, and even Psychological Signals and Audio inputs—a richer, more precise understanding of the driving environment can be achieved. Approaches like multimodal fusion, which integrates data from these varied sources, cross-modality, which allows models to leverage information across different types of data, and shared representation learning, which finds commonalities across these data types, are crucial. This expanded use of multimodal signals promises to enhance the robustness, adaptability, and overall performance of car-following models, allowing them to better predict and react to the dynamic complexities of real-world driving.
Self-Supervised Learning: Traditional car-following models often rely heavily on specific objectives, predefined parameters, and extensively labeled data, which can limit their adaptability and applicability to new or unseen driving scenarios. Self-Supervised Learning (SSL) represents a paradigm shift, enabling models to learn from data that is not explicitly labeled for a specific prediction task. By exploiting the intrinsic structure of driving data, SSL techniques can uncover complex, latent relationships between past and future driving states without the need for direct supervision. This is achieved through mechanisms such as contrastive learning, which distinguishes between similar and dissimilar sequences of driving behavior, and regularized learning, which imposes additional constraints to guide the learning process. Moreover, SSL opens up new avenues for creating models that can adapt over time, learning from continuous streams of data generated by vehicles in operation. This ongoing learning process allows models to evolve in response to changing driving conditions and behaviors, significantly enhancing their long-term utility and effectiveness. The adoption of self-supervised learning in car-following models marks a critical step towards developing autonomous driving systems that can operate with an unprecedented level of sophistication and autonomy, mirroring the complex decision-making processes of human drivers.
VII acknowledgement